OpenStudy (anonymous):

Line cd passes through point (0,2) and (4,6). What is the equation in standard form

OpenStudy (anonymous):

The standard form is: \[y=ax+b\] so all you have to do is insert the 2 coordinates and make 2 equations with 2 variables \[2=a*0+b\]\[6=a*4+b\]

OpenStudy (rhystic):

So you have the definition of a line which is: y = mx + b you can first find m, the slope, by utilizing (y2-y1)/(x2-x1) [slope = rise over run] which yields: (6-2)/(4-0) after you have m you need b. utilizing one set of coordinates: y = m(x) +b sub in the coordinates yielding : 2 = (1)0 + b = > b = 2 you can double check by utilizing the second set of coord : 6=1(4)+b -> b = 2 thus you now have everything for the standard form equation: y = mx + b y = (1)x + 2